An amount borrowed at 29% interest per year, compounded continuously, has accumulated to R38 279,20 after four years. The initial amount borrowed was

[1] R13 823,05.

[2] R12 000,00.

[3] R12 005,53.

[4] R17 721,85.

[5] R7 160,73.


1
Expert's answer
2021-05-24T19:08:48-0400

"Future \space value=Amount\times e^{rt}"

"3827920=Amount\times e^{0.29\times 4}"

"3827920=Amount\times 3.189933"

"\frac{3827920}{3.189933}=Amount"

"Amount=1200000.022"

"Amount=1200000"

[2] R12 000,00. is correct option.